Palencia Paints Corporation has a target capital structure of 40% debt and 60% common equity, with no preferred stock. Its before-tax cost of debt is 10%, and its marginal tax rate is 25%. The current stock price is P0 = $34.50. The last dividend was D0 = $2.50, and it is expected to grow at a 4% constant rate. What is its cost of common equity and its WACC? Do not round intermediate calculations. Round your answers to two decimal places.
Answer #1

P0 = D1/(Ke - g)
D1 = $2.50 * (1+4%) = $2.60
g = 4%
P0 = $34.50

$34.50 = $2.60/(Ke - 0.04)

ke - 0.04 = $2.60/$34.50

Ke - 0.04 = 0.07536

Ke = 0.0754 + 0.04 = 11.54%

Cost of common equity = 11.54%

WACC = (Weight of debt * Pretax debt cost * (1 - Tax)) + (Weight of Equity * Cost of Equity)

= (0.40 * 10% * (1-0.25) + (0.60 * 11.54%)

= 3% + 6.92%

= 9.92%

WACC = 9.92%